OpenStudy (anonymous):

Is 3 - x the same as x + 3 (since you rearrange the terms?)?

OpenStudy (anonymous):

Oops, x - 3

OpenStudy (texaschic101):

3 - x is not the same as x - 3 Because in the first one, the x is negative and the 3 is positive and in the second one, the x is positive and the 3 is negative.

OpenStudy (texaschic101):

now...3 - x is the same as -x + 3

OpenStudy (anonymous):

Thanks!

OpenStudy (texaschic101):

no problem :)

OpenStudy (mathstudent55):

In order for two polynomials to be equal, all terms of one polynomial must equal all corresponding terms of the other polynomial. A polynomial of more than one term is a SUM of terms. Subtraction can be written as the sum of the opposite. For example, 5 - 3 is the same as 5 + (-3). Similarly, x - 6 can be written as x + (-6) which is now a sum. Let's look at both of your polynomials and write them as sums of terms. First polynomial: 3 - x 3 - x = 3 + (-x) In descending order of degree, then 3 - x = 3 + (-x) = -x + 3 Second polynomial: x - 3 x - 3 = x + (-3) The terms of the first polynomial are -x and 3. The terms of the second polynomial are x and -3. Since in general x≠−x and since \(3 \ne -3\), the corresponding terms are not equal, and the polynomials are not equal. Notice what the polynomials are, though. They are opposites, or additive inverses, since they add to zero. (3 - x ) + (x - 3 ) = 3 - x + x - 3 = 3 - 3 - x + x = 0